# HOMIE: Humanoid Loco-Manipulation with Isomorphic Exoskeleton Cockpit

### v1

@@ Appendix B Hardware System Details
In Fig. 1, we present the hardware system design framework of HOMIE, which comprises three integral components: isomorphic exoskeleton arms, a pair of motion-sensing gloves, and a pedal. The primary structural elements of these components are fabricated using 3D printing technology with PLA basic material.
@@ B-B Motion-sensing Glovesc Details
The motion-sensing gloves are secured to the palm via a length-adjustable elastic strap and connected to the fingertips through five smaller length-adjustable elastic straps, which facilitate finger fixation and critical angle mapping, ensuring adaptability to operators with varying hand sizes. A microcontroller is embedded within the palm section of the gloves, featuring exposed ports that allow direct connection to the 15 Hall sensor modules located at the fingers, as depicted in Fig. 17(a). For the joint mapping angle range and acquisition accuracy of each finger, we have listed the data in TABLEx. Each finger of the glove has three degrees of freedom, which are shown in Fig. 17(a), namely the pitch motion of the fingertip ($\alpha$), the pitch motion of the finger pad ($\beta$), and the yaw motion of the finger pad ($\gamma$). Due to differences in the structural length of the thumb, the pinky, and the other three fingers, we have divided them into three parts. It should be noted that the angular movement of the finger joints does not exhibit a significant linear relationship with the changes in the Hall sensor readings caused by the induced magnetic field variation. This is related to the positioning of the magnets and Hall sensors, as well as the structural design of the gloves. In our motion-sensing gloves design, the relationship between the two follows an exponential pattern within their transformation range, especially in the pitch motion of the fingers and finger pads, from open to fist. Furthermore, our gloves have undergone control testing on the Inspire Dexterous Hands RH56DFTP actual device, illustrated in Fig. 17(b).
Figure 17: (a): The schematic diagrams of the three degrees of freedom for each finger and the rotation of the magnet affecting the magnetic field direction, as well as the photograph of the actual wearable device. (b): Physical images of the Inspire Hands actual device in both open and clenched fist states.

### v2

@@ Appendix B Hardware System Details
In Fig. 1, we present the hardware system design framework of HOMIE, which comprises three integral components: isomorphic exoskeleton arms, a pair of motion-sensing gloves, and a pedal. The primary structural elements of these components are fabricated using 3D printing technology with PLA basic material. PLA provides adequate strength at a low cost, and 3D-printed components can be easily built and modified. Building hardware with other materials is feasible using our CAD models.
@@ B-B Motion-sensing Glovesc Details
The Unitree G1 is equipped with the Unitree Dex3-1, a three-fingered dexterous hand with 7 DoF(three in the thumb and two each in the middle and index fingers). Our motion-sensing gloves can track up to 15 DoF, enabling direct mapping of thumb, middle finger, and index finger movements to the corresponding fingers on the Unitree Dex3-1. The motion-sensing gloves are secured to the palm via a length-adjustable elastic strap and connected to the fingertips through five smaller length-adjustable elastic straps, which facilitate finger fixation and critical angle mapping, ensuring adaptability to operators with varying hand sizes. A microcontroller is embedded within the palm section of the gloves, featuring exposed ports that allow direct connection to the 15 Hall sensor modules located at the fingers, as depicted in Fig. 19(a). For the joint mapping angle range and acquisition accuracy of each finger, we have listed the data in Tab. IX. Each finger of the glove has three degrees of freedom, which are shown in Fig. 19(a), namely the pitch motion of the fingertip ($\alpha$), the pitch motion of the finger pad ($\beta$), and the yaw motion of the finger pad ($\gamma$). Due to differences in the structural length of the thumb, the pinky, and the other three fingers, we have divided them into three parts. It should be noted that the angular movement of the finger joints does not exhibit a significant linear relationship with the changes in the Hall sensor readings caused by the induced magnetic field variation. This is related to the positioning of the magnets and Hall sensors, as well as the structural design of the gloves. In our motion-sensing gloves design, the relationship between the two follows an exponential pattern within their transformation range, especially in the pitch motion of the fingers and finger pads, from open to fist. Furthermore, our gloves have undergone control testing on the Inspire Dexterous Hands RH56DFTP actual device, illustrated in Fig. 19(b).
Figure 19: (a): The schematic diagrams of the three degrees of freedom for each finger and the rotation of the magnet affecting the magnetic field direction. (b): Physical images of the Inspire Hands actual device in both open and clenched fist states.
@@ C-C Simulation Deployment
Appendix D User Study
D-A User Study Raw Data
We list raw data got in user study in Tab. X. Our testers represented a diverse spectrum of heights, body types, and genders performing handover tasks. All participants were equipped with identical exoskeleton hardware throughout trials. Despite anthropometric variations from the system’s design specifications, every tester achieved expert-level proficiency within minimal practice time, demonstrating the system’s strong adaptability to diverse body types and inherent user-friendliness.
TABLE X: User Study Raw Data
User ID User Height (cm) User Weight (kg) User Gener Time1(s) Time2(s) Time3(s) Time4(s) Time5(s)
1 185 71 Male 34 28 15 16 9
2 185 90 Male 19 16 13 7 6
3 160 55 Female 19 28 14 17 12
4 170 63 Male 29 21 14 13 6
5 175 80 Male 23 22 15 12 7
Average - - - 24.8 23.0 14.2 13.0 8.0
D-B Teaching Procedure
The standardized onboarding protocol comprises six key phases as shown below, with operational objectives paired with corresponding verbal guidance examples:
  • Customized exoskeleton fitting with adjustable straps for user-centric anthropometric adaptation: “Please wear the exoskeleton backpack-style by fastening the two back straps, then secure the finger straps on each fingertip. Adjust strap tightness according to your comfort level.”
  • Ergonomic positioning featuring seat-height optimization and pedal placement calibration: “Please sit on the chair and adjust the pedal distance to your legs for optimal comfort. Try pressing the pedal to experience the operation.”
  • Systematic briefing on pedal control layout with functional mapping visualization: “The white center pedal controls robot squatting. The large right pedal controls forward/backward movement - press the small right pedal to toggle direction. The large left pedal controls turning, with its small pedal switching left/right orientation.”
  • Motion-sensing glove familiarization through bidirectional mapping demonstrations: “The dexterous hand control program is now active. Move your fingers to experience how each robotic finger’s 2-3 joints correspond to your glove’s knuckle joints and fingertip sensors.”
  • Immersive exoskeleton-arm coordination training via real-time manipulation exercises: “The full-body control program is activated. Move your arms to observe mirrored robot movements. Use pedals for squatting or locomotion commands.”
  • Structured task execution guidance for teleoperation mastery: “You may now combine exoskeleton movements with pedal operations to complete loco-manipulation tasks through the humanoid robot.”
The entire protocol is typically completed within 5 minutes from initial fitting to task execution.